$SPAD/src/input cdraw.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- complex surface and vector field drawing by SCM
-- complex surface vector field drawing
C := Complex DoubleFloat
S := Segment DoubleFloat
PC := Record(rr:DoubleFloat, th:DoubleFloat)
realSteps: PI := 25    -- the number of steps in the real direction
imagSteps: PI := 25    -- the number of steps in the imaginary direction
clipValue: DoubleFloat := 10    -- the maximum length of a vector to draw

-- Draw a complex function as a height field
-- uses the complex norm as the height and the complex argument as the color
-- optionally it will draw arrows on the surface indicating the direction
-- of the complex argument
-- sample call:
--   f: C -> C
--   f z == exp(1/z)
--   drawComplex(f, 0.3..3, 0..2*%pi, false)
-- parameter descriptions:
--   f:  the function to draw
--   rRange: the range of the real values
--   imagRange: the range of imaginary values
drawComplex(f: C -> C, realRange: S, imagRange: S): VIEW3D ==
  free realSteps, imagSteps
  delReal := (hi(realRange) - lo(realRange))/realSteps
  delImag := (hi(imagRange) - lo(imagRange))/imagSteps
  funTable: ARRAY2(PC) := new(realSteps+1, imagSteps+1, [0,0]$PC)
  real := lo(realRange)
  for i in 1..realSteps+1 repeat
    imag := lo(imagRange)
    for j in 1..imagSteps+1 repeat
      z := f complex(real, imag)
      funTable(i,j) := [clipFun(sqrt norm z), argument(z)]$PC
      imag := imag + delImag

  real := real + delReal
llp:List List Point DoubleFloat := []
real := lo(realRange)
for i in 1..realSteps+1 repeat
  imag := lo(imagRange)
  lp:List Point DoubleFloat := []
  for j in 1..imagSteps+1 repeat
    lp := cons(point [real,imag, funTable(i,j).rr,
                                  funTable(i,j).th] ,lp)
    imag := imag + delImag
  real := real + delReal
  llp := cons(reverse! lp, llp)
llp := reverse! llp
space := mesh(llp)$ThreeSpace(DoubleFloat)
makeViewport3D(space, "Complex Function")$VIEW3D
-- draw a complex vector field
-- these vector fields should be viewed from the top by pressing the
-- "XY" translate button on the VIEW3D control panel
-- parameters:
--   f: the mapping from C to C which we will draw
--   realRange: the range of the reals
--   tRange: the range of the imaginaries
-- sample call:
--    f z == sin z
--    drawComplexVectorField(f, -2..2, -2..2)
-- call the functions 'setRealSteps' and 'setImagSteps' to change the
-- number of arrows drawn in each direction.
drawComplexVectorField(f: C -> C, realRange: S, imagRange: S): VIEW3D ==
-- compute the steps size of the grid
delReal := (hi(realRange) - lo(realRange))/realSteps
delImag := (hi(imagRange) - lo(imagRange))/imagSteps
-- create the space to hold the arrows
space := create3Space()$ThreeSpace DoubleFloat
real := lo(realRange)
for i in 1..realSteps+1 repeat
  imag := lo(imagRange)
  for j in 1..imagSteps+1 repeat
    -- compute the function
    z := f complex(real, imag)
    -- get the direction of the arrow
    arg := argument z
    -- get the length of the arrow
    len := clipFun(sqrt norm z)
    -- create point at the base of the arrow
    p1 :=  point [real, imag, 0.0@DoubleFloat, arg]
    -- scale the arrow length so it isn't too long

      scaleLen := delReal * len
      -- create the point at the top of the arrow
      p2 := point [p1.1 + scaleLen*cos(arg), p1.2 + scaleLen*sin(arg),
                   0.0@DoubleFloat, arg]
      -- make the pointer at the top of the arrow
      arrow := makeArrow(p1, p2, scaleLen, arg)
      -- add the line segments in the arrow to the space
      for a in arrow repeat curve(space, a)$ThreeSpace DoubleFloat
      imag := imag + delImag
    real := real + delReal
  -- draw the vector feild
  makeViewport3D(space, "Complex Vector Field")$VIEW3D
-- relative size of the arrow head compared to the length of the arrow
arrowScale := 0.25@DoubleFloat
-- angle of the arrow head
arrowAngle := %pi-%pi/10.0@DoubleFloat
-- Add an arrow head to a line segment, which starts at 'p1', ends at 'p2',
-- has length 'len', and and angle 'arg'.  We pass 'len' and 'arg' as
-- arguments since thet were already computed by the calling program
makeArrow(p1, p2, len, arg) ==
  c1 := cos(arg + arrowAngle)
  s1 := sin(arg + arrowAngle)
  c2 := cos(arg - arrowAngle)
  s2 := sin(arg - arrowAngle)
  p3 := point [p2.1 + c1*arrowScale*len, p2.2 + s1*arrowScale*len,
               p2.3, p2.4]
  p4 := point [p2.1 + c2*arrowScale*len, p2.2 + s2*arrowScale*len,
               p2.3, p2.4]
  [[p1, p2, p3], [p2, p4]]
-- set the number of steps to use in the real direction
setRealSteps(n) ==
  free realSteps
  realSteps := n
-- set the number of steps to use in the imaginary direction
setImagSteps(n) ==
  free imagSteps
  imagSteps := n
-- set the maximum length of a vector
setClipValue clip ==
  free clipValue
  clipValue := clip
-- clip a value in the interval (-clip...clip)
clipFun(x:DoubleFloat):DoubleFloat ==

min(max(x, -clipValue), clipValue)

———-

References^{[1] nothing}

7